OpenStudy (anonymous):

Triangle END is translated using the rule (x,y)→(x-4, y-1) to create triangle E'N'D'. If a line segment is drawn from point E to point E' and from point N to point N', which statement would best describe the line segments drawn? Select one: a. They are parallel and congruent. b. They are perpendicular to each other. c. They share the same midpoints. d. They create diameters of concentric circles.

OpenStudy (anonymous):

i really need help

OpenStudy (anonymous):

hey

ganeshie8 (ganeshie8):

give some coordinates, and try to find slope of EE' and NN'

OpenStudy (anonymous):

wat coordinates am i suppossed to show

ganeshie8 (ganeshie8):

say, \(E=(x_1,y_1)\) and, \(N=(x_2,y_2) \) after (x,y)-->(x-4, y-1) translation, \(E'=(x_1-4,y_1-1)\) and, \(N'=(x_2-4,y_2-1) \)

ganeshie8 (ganeshie8):

now, find slope of EE' and NN'

OpenStudy (anonymous):

ok ill try

ganeshie8 (ganeshie8):

ok give it a try, use the slope formula

OpenStudy (anonymous):

am i supposed to be plugging in numbers

ganeshie8 (ganeshie8):

plugin the coordinates

OpenStudy (anonymous):

wat are the coordinates

ganeshie8 (ganeshie8):

slope of \(\large E(x_1, y_1)\) and \(E'(x_1-4, y_1-1) = \frac{y_1-1-y_1}{x_1-4-x_1}\)

ganeshie8 (ganeshie8):

simplify

OpenStudy (anonymous):

man i dont get this.... :(

ganeshie8 (ganeshie8):

np, we're trying to do this :- u simplify it. after that, find slope of NN', and simplify it aswell. you will get same value for both. then, u can conclude that, the lines are parallel

ganeshie8 (ganeshie8):

btw, answer is q, when u do the listed steps above, u wil see that EE' & NN' are parallel and and congruent

OpenStudy (anonymous):

so they will be parrallel and congruent right

ganeshie8 (ganeshie8):

yes *answer is a

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

lets do another